The length of a bicycle pedal arm is 0.157 m, and a downward force of 127 N is applied to the pedal by the rider. What is the magnitude of the torque about the pedal arm's pivot when the arm is at an angle of (a)18.6°, (b) 90°, and (c) 180° with the vertical?
Answer #1

F = 127N , r = 0.157m

Since torque T = rF Sin(theta)

A.) Theta = 18.6°

T = (0.157)(127) Sin(18.6) = 6.36N-m

=) T = 6.36N-m

B.) Theta = 90°

T = (0.157)(127)(1) = 19.94 N-m

C.) Theta = 180°

T = (0.157)(127) Sin(180°) = 0 .